OpenStudy (anonymous):

can anyone break down the steps in this problem, so I can understand it. Solve for x. log24 – log4x = 0 a.54 b.16 c. 6 d.36

OpenStudy (anonymous):

Do either of these have abnormal bases?

OpenStudy (anonymous):

do you mean the 24?

OpenStudy (anonymous):

yeah, is it log base 2, log base 24, or is it just log(24)-log(4x)?

OpenStudy (anonymous):

it is the third one you wrote

OpenStudy (anonymous):

okay, give me just a sec to write this out.

OpenStudy (anonymous):

So what you're going to do is cancel out the log part of the expression. To do this, make each part of the equation an exponent to the number 10. You make it an exponent of 10 because log has a base of 10. Like this: \[10^{\log(24)}-10^{\log(4x)}\] This makes the equation a simple 24-4x=0 and then you just solve for x. \[-4x=-24\]\[x=-24\div-4\]\[x=6\] Does this answer your question?

OpenStudy (anonymous):

yes thank you so much.. I have been following my lessons and I just got lost this really helped (:

OpenStudy (anonymous):

You're welcome :)